# Feature Selection for Functional Data Classification

### 11 Jan 2024 v1

@@ Abstract
Functional data analysis has emerged as a crucial tool in many contemporary scientific domains that require the integration and interpretation of complex data. Moreover, the advent of new technologies has facilitated the collection of a large number of longitudinal variables, making feature selection pivotal for avoiding overfitting and improving prediction performance. This paper introduces a novel methodology, called FSFC (Feature Selection for Functional Classification), that addresses the challenge of jointly performing feature selection and classification of functional data in scenarios with categorical responses and longitudinal features. Our approach tackles a newly defined optimization problem that integrates logistic loss and functional features to identify the most crucial features for classification. To address the minimization procedure, we employ functional principal components and develop a new adaptive version of the Dual Augmented Lagrangian algorithm that leverages the sparsity structure of the problem for dimensionality reduction. The computational efficiency of FSFC enables handling high-dimensional scenarios where the number of features may considerably exceed the number of statistical units. Simulation experiments demonstrate that FSFC outperforms other machine learning and deep learning methods in computational time and classification accuracy. Furthermore, the FSFC feature selection capability can be leveraged to significantly reduce the problem’s dimensionality and enhance the performances of other classification algorithms. The efficacy of FSFC is also demonstrated through a real data application, analyzing relationships between four chronic diseases and other health and socio-demographic factors.
Introduction
Feature selection is among the most renowned and widely utilized techniques in machine learning (ML) and deep learning (DL) to decrease the dimensionality of optimization problems, by identifying a subset of essential features, and it has been applied to different tasks including linear regression (Tibshirani 1996; Zou and Hastie 2005) and classification (Cai et al. 2018). Moreover, in many modern applications, the number of features is much larger than the number of statistical units (Ovsyannikova et al. 2020; Paul et al. 2014), and in such cases reducing the problem dimensionality is crucial since ML and DL models may risk overfitting and lose efficiency (Guyon et al. 2002; Zhang et al. 2021).
@@ Introduction
Instead, Functional Data Analysis (FDA) (Ramsay and Silverman 2005; Horváth and Kokoszka 2012) is an active domain within statistics and ML that allows to work with longitudinal measurements by estimating smooth curves from multiple observations over a continuous domain. FDA has demonstrated its efficacy in many recent applications where the emergence of novel technologies, such as brain sensors, DNA sequencers, and wearable devices, allows to measure various signals over time (Shcherbina et al. 2017; Cremona et al. 2019; Smuck et al. 2021).
This paper introduces a new methodology called Feature Selection for Functional Classification (FSFC) to perform feature selection with functional data for a classification task, i.e., in scenarios where features are longitudinal, and responses are categorical. There are numerous existing methods for classifying longitudinal data, including supporting vector machines (SVM) (Müller et al. 1997), long short-term memory (LSTM) networks (Hochreiter and Schmidhuber 1997), shapelets (Ye and Keogh 2009), and functional classification (Leng and Müller 2006; Fraiman, Gimenez, and Svarc 2016). Additionally, feature selection in the context of linear functional regression has been extensively studied (Chen, Goldsmith, and Ogden 2016; Parodi, Reimherr et al. 2018; Boschi, Reimherr, and Chiaromonte 2021). However, to the best of our knowledge, no study to date has been conducted on approaches that enable simultaneous feature selection and classification within the framework of functional data.
The main contributions of this work can be summarized as follows:
@@
    We develop a novel methodology, FSFC, designed to concurrently perform feature selection and classification of functional data in scenarios with categorical responses and longitudinal features. Specifically, FSFC solves a newly defined optimization function, identifying the most important features for the classification task.
@@
    We employ functional principal components and introduce an innovative adaptive version of the Dual Augmented Lagrangian (DAL) algorithm to address the optimization problem. The computational efficiency of FSFC enables tackling high-dimensional scenarios, where the number of features may considerably exceed the number of statistical units.
@@
    We investigate the efficacy of FSFC and its potential to enhance other ML and DL methods through simulations and a real data application, where we examine the relationships between four chronic diseases (namely, diabetes, myocardial infarction, high cholesterol, and hypertension) and other health and socio-demographic factors.
In the remainder of this paper, we first present our novel methodology, and then we benchmark its performance against SVMs and LSTMs via simulations and a real-world application on the longitudinal SHARE dataset (Alcser et al. 2005).
Methods
Problem definition
@@ Methods
The first term of the objective function (1) is the logistic loss with functional features and enables classification. The second term comprises two different penalties that induce sparsity to facilitate feature selection. The parameters $\lambda_{1}$ and $\lambda_{2}$ control the global importance of the penalties with respect to the logistic loss. The feature-specific weights $\omega_{j}$ extend the adaptive LASSO and Elastic Net (Zou 2006; Zou and Zhang 2009) to the functional classification settings and can improve feature selection by reducing the active set cardinality.
In order to solve (1), first, we approximate the functional variables via the Functional Principal Components (FPC) method (see, e.g. Horváth and Kokoszka 2012), and subsequently, perform the minimization process using a Dual Augmented Lagrangian (DAL) algorithm. The DAL algorithm was introduced Tomioka and Sugiyama (2009); Li, Sun, and Toh (2018) to solve LASSO and Elastic-Net problems and has been extended by Boschi et al. (2023) to accommodate functional linear regression cases. DAL leverages the intrinsic sparsity of the problem and the sparsity induced by the Hessian matrix information, resulting in a substantial reduction in computational cost. In this work, we devise a novel adaptive version of DAL and develop the mathematical theory underpinning the algorithm to incorporate logistic loss and functional features within the objective function.
Matrix representation
To address (1), we employ a matrix representation obtained by expressing functional variables as linear combinations of basis functions (Ramsay and Silverman 2005; Kokoszka and Reimherr 2017). Specifically, we represent the functional features using their first $k$ FPCs. FPCs are a well-established technique to solve the functional regression (Reiss and Ogden 2007) and classification (Preda, Saporta, and Lévéder 2007; Wang, Chiou, and Müller 2016) problems. We now show how to use FPCs to obtain a matrix representation of (1).
@@
$\begin{split}\min_{B}~{}\Bigg{[}\sum_{i=1}^{n}\log\bigg{(}1+\exp\Big{(}-Y_{i}\cdot\big{(}X_{(i)}B\big{)}\Big{)}\bigg{)}\\ +\sum_{j=1}^{p}\omega_{j}\left(\lambda_{1}\lVert B_{j}\rVert_{2}+\frac{\lambda_{2}}{2}\lVert B_{j}\rVert_{2}^{2}\right)\Bigg{]}\ ,\end{split}$ (2)
where $X_{(i)}$ indicates the $i$-th row of $X$. Similar to Elastic Net regularization (Zou and Hastie 2005), Equation (2) combines two penalties: the first is non-differentiable, creating sparsity, while the second (Ridge-like) is differentiable, controlling multicollinearity and accelerating the convergence of the optimization algorithm. Once we obtain an estimate of the score vector $\hat{B}$, the coefficient curves can be recovered as ^$\hat{\mathcal{B}}_{j}(t)=e^{j}(t)\hat{B}_{j}(t)$, the class probability as $\hat{p}_{i}=1/\big{(}1+\exp\big{(}X_{(i)}\hat{B}\big{)}\big{)}$, and the categorical response as $\hat{Y}_{i}=1$ if $\hat{p}_{i}>0.5$ and $-1$ otherwise.
@@
Dual Augmented Lagrangian (DAL) Algorithm
@@
A potential dual representation of the primal (P) can be derived from (Boyd and Vandenberghe 2004; Tomioka, Suzuki, and Sugiyama 2011) as:
@@
In this formulation, $V\in\mathbb{R}^{n}$ and $Z\in\mathbb{R}^{pk}$ are the dual variables. Following the notation introduced for $B$, we can express $Z=\left(Z_{1}^{T},\dots,Z_{p}^{T}\right)^{T}$, where $Z_{j}\in\mathbb{R}^{k}$ denotes the sub-vector of $Z$ corresponding to the $j$-th feature. $h^{*}$ and $\pi^{*}$ denote the Fenchel-conjugate functions (Fenchel 1949) of $h$ and $\pi$, respectively. The computation of $\pi^{*}$ is derived in (Boschi et al. 2023) as
$\begin{split}\pi^{*}(Z)=\sum_{j=1}^{p}\pi^{*}(Z_{j})=\sum_{j=1}^{p}(2\omega_{j}\lambda_{2})^{-1}\left(\big{[}\left\lVert Z_{j}\right\rVert_{2}-\omega_{j}\lambda_{1}\big{]}_{+}\right)^{2},\end{split}$ (3)
@@
$\begin{split}h^{*}(V)=\sum_{i=1}^{n}(1-\left|Y_{i}V_{i}\right|)\log(1-\left|Y_{i}V_{i}\right|)+\left|Y_{i}V_{i}\right|\log(\left|Y_{i}V_{i}\right|)\ .\end{split}$ (4)
@@
($\begin{split}\mathcal{L}&{}_{\sigma}(V,Z,B)=h^{*}(V)+\pi^{*}(Z)\\ &-\sum_{j=1}^{p}\langle B_{j},V^{T}X_{j}+Z_{j}\rangle+\frac{\sigma}{2}\sum_{j=1}^{p}\lVert V^{T}X_{j}+Z_{j}\rVert_{2}^{2}\ ,\end{split}$ (5)
@@
Our DAL implementation is outlined in Algorithm 1, as adapted from Boschi et al. (2023). The core part of the algorithm involves solving the inner sub-problem (6), which entails minimizing the Dual Augmented Lagrangian with respect to the dual variables $V$ and $Z$. To obtain an approximate solution of (6), $V$ and $Z$ can be updated independently (Tomioka and Sugiyama 2009). The primal variable $B$ and the parameter $\sigma$ are updated according to standard rules (Li, Sun, and Toh 2018). The choice of $\sigma^{0}$ is empirically studied in Tomioka, Suzuki, and Sugiyama (2011): starting from very small values of $\sigma$ results in an increased number of iterations required for convergence; however, if $\sigma^{0}$ is excessively large or increases too rapidly, the DAL method fails to converge to the optimal solution.
@@
The minimization with respect to $Z$ has a closed-form solution $\bar{Z}$, which is computed in Boschi et al. (2023) based on the proximal operator (Rockafellar 1976a, b) of $\pi$, denoted as $\operatorname{prox}_{\sigma\pi}$. Specifically, the following holds:
$\bar{Z}$ $=\operatorname{prox}_{\pi^{*}/\sigma}\left(B/\sigma-X^{T}\bar{V}\right)$
@@
$\operatorname{prox}_{\sigma\pi}(B_{j})=(1+\sigma\omega_{j}\lambda_{2})^{-1}\left[1-\left\lVert B_{j}\right\rVert_{2}^{-1}\sigma\omega_{j}\lambda_{1}\right]_{+}B_{j}\ .$ (8)
@@
Given that the minimization with respect to $V$ lacks a closed-form solution, we update $V$ applying the Newton Method (Nocedal and Wright 1999) to the function $\psi(V):=\mathcal{L}_{\sigma}(V\,|\,\bar{Z},B)$. The Newton update is expressed by $V^{m+1}=V^{m}+sD$, where $D\in\mathbb{R}^{n}$ represents the descent direction and $s$ the step-size. To obtain $D$, we solve the linear system
@@
where $H_{\psi}\in\mathbb{R}^{n\times n}$ and $\nabla\psi\in\mathbb{R}^{n}$ are the Hessian matrix and the gradient vector of $\psi$, respectively. To determine the step size, we implement the line-search strategy proposed by Li, Sun, and Toh (2018). Starting with $s=1$, we iteratively reduce it by a factor of $0.5$ until the condition
@@
$\begin{split}P_{[j]}=(1+\sigma\omega_{j}\lambda_{2})^{-1}\left(\left(1-\frac{\sigma\omega_{j}\lambda_{1}}{\left\lVert T_{j}\right\rVert_{2}}\right)I_{k}+\frac{\sigma\omega_{j}\lambda_{1}}{\left\lVert T_{j}\right\rVert_{2}^{3}}T_{j}T_{j}^{T}\right)\ .\end{split}$
@@
$\begin{split}&(i)~{}\ \ \psi(V)=h^{*}(V)\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\frac{1}{2\sigma}\sum_{j=1}^{p}\left(\left(1+\sigma\omega_{j}\lambda_{2}\right)\big{\lVert}\operatorname{prox}_{\sigma\pi}\big{(}T_{j}\big{)}\big{\rVert}_{2}^{2}-\left\lVert B_{j}\right\rVert_{2}^{2}\right)\ ,\\ &(ii)~{}\ \nabla\psi(V)=\nabla h^{*}(V)-X\operatorname{prox}_{\sigma\pi}(T)\ ,\\ &(iii)~{}H_{\psi}(V)=H_{h^{*}}(V)+\sigma X_{\mathcal{J}}Q_{\mathcal{J}}X_{\mathcal{J}}^{T}\ ,\end{split}$
@@
As stated in Proposition 2, the DAL algorithm takes advantage of the sparsity information embedded within the Hessian matrix, selecting a subset $\mathcal{J}$ of $r$ active features, with $r$ decreasing at each iteration. By selecting an $r$-feature subset, the overall computational expense of solving the linear system (9) is reduced from $\mathcal{O}\left(n(n^{2}+npk+p^{2}k^{2})\right)$ to $\mathcal{O}\left(n(n^{2}+nrk+r^{2}k^{2})\right)$ with $r<p$. In sparse scenarios, where the number of active features is low, $r$ might be significantly smaller than $n$, further reducing the computational burden through the application of the Sherman-Morrison-Woodbury formula (Van Loan and Golub 1983): $\left(H_{h^{*}}+\sigma X_{\mathcal{J}}Q_{\mathcal{J}}X_{\mathcal{J}}^{T}\right)^{-1}=H_{h^{*}}^{-1}-H_{h^{*}}^{-1}X_{\mathcal{J}}\big{(}\left(\sigma Q_{\mathcal{J}}\right)^{-1}+X_{\mathcal{J}}^{T}H_{h^{*}}^{-1}X_{\mathcal{J}}\big{)}^{-1}X_{\mathcal{J}}^{T}H_{h^{*}}^{-1}$. This equivalence enables the factorization of an $rk\times rk$ matrix, resulting in a total cost of $\mathcal{O}\left(rk(k^{2}+nrk+r^{2}k^{2}+n^{2})\right)$ since $H_{h^{*}}$ is diagonal and computing its inverse is straightforward. Remarkably, the computational burden is not dependent on the total number of features $p$, but solely on the number of active features $r$. Provided that sparsity is maintained (i.e., $r$ remains small), the number of features can grow without impacting the efficiency of the linear system resolution. However, it is important to acknowledge that a larger $p$ increases the cost of the principal components computation during the matrix representation stage, since the FPC scores have to be derived for each feature. Lastly, note that the computational cost depends on $k^{3}$. By maintaining a small value for $k$, which describes the number of bases used to represent the functional features, the efficiency of the method is significantly enhanced.

### 05 Mar 2024 v2

# A new computationally efficient algorithm to solve Feature Selection for Functional Data Classification in high-dimensional spaces

@@ Abstract
This paper introduces a novel methodology for Feature Selection for Functional Classification, FSFC, that addresses the challenge of jointly performing feature selection and classification of functional data in scenarios with categorical responses and multivariate longitudinal features. FSFC tackles a newly defined optimization problem that integrates logistic loss and functional features to identify the most crucial variables for classification. To address the minimization procedure, we employ functional principal components and develop a new adaptive version of the Dual Augmented Lagrangian algorithm. The computational efficiency of FSFC enables handling high-dimensional scenarios where the number of features may considerably exceed the number of statistical units. Simulation experiments demonstrate that FSFC outperforms other machine learning and deep learning methods in computational time and classification accuracy. Furthermore, the FSFC feature selection capability can be leveraged to significantly reduce the problem’s dimensionality and enhance the performances of other classification algorithms. The efficacy of FSFC is also demonstrated through a real data application, analyzing relationships between four chronic diseases and other health and demographic factors.
Tobia Boschi1, *, Francesca Bonin1, Rodrigo Ordonez-Hurtado1,
Alessandra Pascale1, and Jonathan Epperlein1
1IBM Research Europe, Dublin
* Corresponding Author: tobia.boschi@ibm.com
1 Introduction
@@
Many contemporary scientific domains that require the integration and interpretation of time-dependent variables have recently benefited from Functional Data Analysis (FDA) (Ramsay & Silverman, 2005; Horváth & Kokoszka, 2012), an active domain within statistics and Machine Learning (ML) that allows working with multivariate longitudinal measurements by estimating smooth curves from multiple observations over a continuous domain. This functional representation facilitates the comparison of different statistical units and variables throughout the entire temporal domain. FDA has demonstrated its effectiveness in many recent applications where the emergence of novel technologies, such as brain sensors, DNA sequencers, and wearable devices, has facilitated the measurement of a large number of longitudinal variables (Shcherbina et al., 2017; Cremona et al., 2019; Smuck et al., 2021). In these modern domain specific applications of FDA, two criticality emerge: 1) high dimensionality, the number of features may significantly exceed the number of statistical units (Ovsyannikova et al., 2020; Paul et al., 2014), 2) data scarcity (Brigato & Iocchi, 2021). The first issue may lead ML and Deep Learning (DL) models to overfit, the second may lead current feature selection models to sub-optimal efficiency.
Hence, there is a growing demand for new algorithms able to i) effectively reduce problem dimensionality and feature complexity, ii) deal with data scarcity (Zhu et al., 2023), and iii) preserve computational efficiency (Brigato & Iocchi, 2021).
To address these challenges in instances characterized by multivariate longitudinal variables and binary categorical responses, we introduce a new feature selection methodology called Feature Selection for Functional Classification (FSFC). The key innovative aspects of our work are the following:
@@
    We define a new optimization problem that integrates logistic loss and multivariate functional features to concurrently solve feature selection and classification.
@@
    We employ Functional Principal Components (FPC) and introduce an innovative adaptive version of the Dual Augmented Lagrangian (DAL) algorithm to address the optimization problem. FPC are used to simplify the complexity of each feature while retaining its longitudinal representation. On the other hand, DAL employs the sparsity structure of the problem to reduce its dimensionality. This combination of FPC and DAL makes FSFC highly computationally efficient, tackling high-dimensional scenarios where the number of features is much larger than the number of statistical units.
@@
    We evaluate the efficacy of FSFC both in isolation and as a pre-processing step to enhance other ML and DL methods; we validate our approach through simulations and a real data application. In the real data application, we examine the relationships between four chronic diseases and other health and socio-demographic factors, demonstrating FSFC adaptability to multi-modal data scenarios.
Related Work.
Feature selection has been a fertile field of research and development since the 1970s and is now among the most renowned and widely utilized techniques in ML and DL to decrease the dimensionality of optimization problems, remove irrelevant and redundant features, improve efficiency in learning and predictive tasks, and enhance model interpretability (Blum & Langley, 1997; Dash & Liu, 1997). It has been applied to several tasks, including linear regression (Tibshirani, 1996; Zou & Hastie, 2005), and classification (Cai et al., 2018) for handling a vast number of instances and when dealing with high-dimensional data (Liu et al., 2002; Das, 2001). Additionally, feature selection has been extensively studied also in the context of linear functional regression (Chen et al., 2016; Parodi et al., 2018; Boschi et al., 2021). However, while there are numerous existing methods for classifying longitudinal observations, including supporting vector machines (SVM) (Müller et al., 1997), long short-term memory (LSTM) networks (Hochreiter & Schmidhuber, 1997), shapelets (Ye & Keogh, 2009), and functional classification (Leng & Müller, 2006; Fraiman et al., 2016), to the best of our knowledge, no study to date has been conducted on approaches that enable simultaneous feature selection and classification within the framework of multivariate functional data. Indeed, although established techniques like Variational AutoEncoders (Kingma et al., 2019) and Time Series Feature Extraction (Barandas et al., 2020) can reduce the complexity of each longitudinal feature by identifying key attributes, they are not designed to decrease the total number of features detecting the most relevant ones and do not preserve their longitudinal information.
In the remainder of this paper, we first present our methodology in Section 2, we benchmark its performance against SVMs and LSTMs via simulations (Section 3) and a real-world application on the longitudinal SHARE dataset (Section 4). We draw our conclusions in Section 5.
2 Methods
2.1 Problem definition
@@
The first term of the objective function (1) is the logistic loss with functional features and enables classification. The second term comprises two different penalties that induce sparsity to facilitate feature selection. The parameters $\lambda_{1}$ and $\lambda_{2}$ control the global importance of the penalties with respect to the logistic loss. The feature-specific weights $\omega_{j}$ extend the adaptive LASSO and Elastic Net (Zou, 2006; Zou & Zhang, 2009) to the functional classification settings and can improve feature selection by reducing the active set cardinality.
In order to solve (1), first, we approximate the functional variables via the Functional Principal Components (FPC) method (see, e.g. Horváth & Kokoszka, 2012), and subsequently, perform the minimization process using a Dual Augmented Lagrangian (DAL) algorithm. The DAL algorithm was introduced Tomioka & Sugiyama (2009); Li et al. (2018) to solve LASSO and Elastic-Net problems and has been extended by Boschi et al. (2023) to accommodate functional linear regression cases. DAL leverages the intrinsic sparsity of the problem and the sparsity induced by the Hessian matrix information, resulting in a substantial reduction in computational cost. In this work, we devise a novel adaptive version of DAL and develop the mathematical theory underpinning the algorithm to incorporate logistic loss and functional features within the objective function.
2.2 Matrix representation
To address (1), we employ a matrix representation obtained by expressing functional variables as linear combinations of basis functions (Ramsay & Silverman, 2005; Kokoszka & Reimherr, 2017). Specifically, we represent the functional features using their first $k$ FPCs. FPCs are a well-established technique to solve the functional regression (Reiss & Ogden, 2007) and classification (Preda et al., 2007; Wang et al., 2016) problems. We now show how to use FPCs to obtain a matrix representation of (1).
@@ 2.2 Matrix representation
$\small\begin{split}\min_{B}~{}\Bigg{[}\sum_{i=1}^{n}\log\bigg{(}1+\exp\Big{(}-Y_{i}\cdot\big{(}X_{(i)}B\big{)}\Big{)}\bigg{)}\\ +\sum_{j=1}^{p}\omega_{j}\left(\lambda_{1}\lVert B_{j}\rVert_{2}+\frac{\lambda_{2}}{2}\lVert B_{j}\rVert_{2}^{2}\right)\Bigg{]}\ ,\end{split}$ (2)
where $X_{(i)}$ indicates the $i$-th row of $X$. Mimicking the Group Elastic Net regularization (Zou & Hastie, 2005; Simon et al., 2013), Equation (2) combines two penalties: the first is non-differentiable, creating sparsity, while the second (Ridge-like) is differentiable, controlling multicollinearity and accelerating the convergence of the optimization algorithm. Once we obtain an estimate of the score vector $\hat{B}$, the coefficient curves can be recovered as ^$\hat{\mathcal{B}}_{j}(t)=e^{j}(t)\hat{B}_{j}(t)$, the class probability as $\hat{p}_{i}=1/\big{(}1+\exp\big{(}X_{(i)}\hat{B}\big{)}\big{)}$, and the categorical response as $\hat{Y}_{i}=1$ if $\hat{p}_{i}>0.5$ and $-1$ otherwise.
@@
2.3 Dual Augmented Lagrangian (DAL) Algorithm
@@
A potential dual representation of the primal (P) can be derived from (Boyd & Vandenberghe, 2004; Tomioka et al., 2011) as:
@@
In this formulation, $V\in\mathbb{R}^{n}$ and $Z\in\mathbb{R}^{pk}$ are the dual variables. Following the notation introduced for $B$, we can express $Z=\left(Z_{1}^{T},\dots,Z_{p}^{T}\right)^{T}$, where $Z_{j}\in\mathbb{R}^{k}$ denotes the sub-vector of $Z$ corresponding to the $j$-th feature. $h^{*}$ and $\pi^{*}$ denote the Fenchel-conjugate functions (Fenchel, 1949) of $h$ and $\pi$, respectively. The computation of $\pi^{*}$ is derived in (Boschi et al., 2023) as
$\small\begin{split}\pi^{*}(Z)=\sum_{j=1}^{p}\pi^{*}(Z_{j})=\sum_{j=1}^{p}(2\omega_{j}\lambda_{2})^{-1}\left(\big{[}\left\lVert Z_{j}\right\rVert_{2}-\omega_{j}\lambda_{1}\big{]}_{+}\right)^{2},\end{split}$ (3)
@@
$\small h^{*}(V)=\sum_{i=1}^{n}(1-\left|Y_{i}V_{i}\right|)\log(1-\left|Y_{i}V_{i}\right|)+\left|Y_{i}V_{i}\right|\log(\left|Y_{i}V_{i}\right|)\ .$ (4)
@@
($\small\begin{split}\mathcal{L}&{}_{\sigma}(V,Z,B)=h^{*}(V)+\pi^{*}(Z)\\ &-\sum_{j=1}^{p}\langle B_{j},V^{T}X_{j}+Z_{j}\rangle+\frac{\sigma}{2}\sum_{j=1}^{p}\lVert V^{T}X_{j}+Z_{j}\rVert_{2}^{2}\ ,\end{split}$ (5)
@@
Our DAL implementation is outlined in Algorithm 1, as adapted from Boschi et al. (2023). The core part of the algorithm involves solving the inner sub-problem (6), which entails minimizing the Dual Augmented Lagrangian with respect to the dual variables $V$ and $Z$. To obtain an approximate solution of (6), $V$ and $Z$ can be updated independently (Tomioka & Sugiyama, 2009). The primal variable $B$ and the parameter $\sigma$ are updated according to standard rules (Li et al., 2018). The choice of $\sigma^{0}$ is empirically studied in Tomioka et al. (2011): starting from very small values of $\sigma$ results in an increased number of iterations required for convergence; however, if $\sigma^{0}$ is excessively large or increases too rapidly, the DAL method fails to converge to the optimal solution.
@@
The minimization with respect to $Z$ has a closed-form solution $\bar{Z}$, which is computed in Boschi et al. (2023) based on the proximal operator (Rockafellar, 1976b, a) of $\pi$, denoted as $\operatorname{prox}_{\sigma\pi}$. Specifically, the following holds:
$\small\bar{Z}$ $=\operatorname{prox}_{\pi^{*}/\sigma}\left(B/\sigma-X^{T}\bar{V}\right)$
@@
$\small\operatorname{prox}_{\sigma\pi}(B_{j})=(1+\sigma\omega_{j}\lambda_{2})^{-1}\left[1-\left\lVert B_{j}\right\rVert_{2}^{-1}\sigma\omega_{j}\lambda_{1}\right]_{+}B_{j}\ .$ (8)
@@
Given that the minimization with respect to $V$ lacks a closed-form solution, we update $V$ applying the Newton Method (Nocedal & Wright, 1999) to the function $\psi(V):=\mathcal{L}_{\sigma}(V\,|\,\bar{Z},B)$. The Newton update is expressed by $V^{m+1}=V^{m}+sD$, where $D\in\mathbb{R}^{n}$ represents the descent direction and $s$ the step-size. To obtain $D$, we solve the linear system
@@
where $H_{\psi}\in\mathbb{R}^{n\times n}$ and $\nabla\psi\in\mathbb{R}^{n}$ are the Hessian matrix and the gradient vector of $\psi$, respectively. To determine the step size, we implement the line-search strategy proposed by Li et al. (2018). Starting with $s=1$, we iteratively reduce it by a factor of $0.5$ until the condition
@@
$\small\begin{split}P_{[j]}=(1+\sigma\omega_{j}\lambda_{2})^{-1}\left(\left(1-\frac{\sigma\omega_{j}\lambda_{1}}{\left\lVert T_{j}\right\rVert_{2}}\right)I_{k}+\frac{\sigma\omega_{j}\lambda_{1}}{\left\lVert T_{j}\right\rVert_{2}^{3}}T_{j}T_{j}^{T}\right)\ .\end{split}$
@@
$\small\begin{split}&(i)~{}\ \ \psi(V)=h^{*}(V)\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\frac{1}{2\sigma}\sum_{j=1}^{p}\left(\left(1+\sigma\omega_{j}\lambda_{2}\right)\big{\lVert}\operatorname{prox}_{\sigma\pi}\big{(}T_{j}\big{)}\big{\rVert}_{2}^{2}-\left\lVert B_{j}\right\rVert_{2}^{2}\right)\ ,\\ &(ii)~{}\ \nabla\psi(V)=\nabla h^{*}(V)-X\operatorname{prox}_{\sigma\pi}(T)\ ,\\ &(iii)~{}H_{\psi}(V)=H_{h^{*}}(V)+\sigma X_{\mathcal{J}}Q_{\mathcal{J}}X_{\mathcal{J}}^{T}\ ,\end{split}$
@@
As stated in Proposition 2, the DAL algorithm takes advantage of the sparsity information embedded within the Hessian matrix, selecting a subset $\mathcal{J}$ of $r$ active features, with $r$ decreasing at each iteration. By selecting an $r$-feature subset, the overall computational expense of solving the linear system (9) is reduced from $\mathcal{O}\left(n(n^{2}+npk+p^{2}k^{2})\right)$ to $\mathcal{O}\left(n(n^{2}+nrk+r^{2}k^{2})\right)$ with $r<p$. In sparse scenarios, where the number of active features is low, $r$ might be significantly smaller than $n$, further reducing the computational burden through the application of the Sherman-Morrison-Woodbury formula (Van Loan & Golub, 1983): $\left(H_{h^{*}}+\sigma X_{\mathcal{J}}Q_{\mathcal{J}}X_{\mathcal{J}}^{T}\right)^{-1}=H_{h^{*}}^{-1}-H_{h^{*}}^{-1}X_{\mathcal{J}}\big{(}\left(\sigma Q_{\mathcal{J}}\right)^{-1}+X_{\mathcal{J}}^{T}H_{h^{*}}^{-1}X_{\mathcal{J}}\big{)}^{-1}X_{\mathcal{J}}^{T}H_{h^{*}}^{-1}$. This equivalence enables the factorization of an $rk\times rk$ matrix, resulting in a total cost of $\mathcal{O}\left(rk(k^{2}+nrk+r^{2}k^{2}+n^{2})\right)$ since $H_{h^{*}}$ is diagonal and computing its inverse is straightforward. Remarkably, the computational burden is not dependent on the total number of features $p$, but solely on the number of active features $r$. Provided that sparsity is maintained (i.e., $r$ remains small), the number of features can grow without impacting the efficiency of the linear system resolution. However, it is important to acknowledge that a larger $p$ increases the cost of the principal components computation during the matrix representation stage, since the FPC scores have to be derived for each feature. Note that the computational cost depends on $k^{3}$. By maintaining a small value for $k$, which describes the number of bases used to represent the functional features, the efficiency of the method is significantly enhanced.